Grade 12th passMechanics

An object starts from rest with constant acceleration 4 m/s^2, then find the distance travelled by object in 5th half seconds.

Vikash Kumar Singh
8 Years agoGrade 12th pass

2 Answers

sonika p
8 Years ago
initial velocity, u=0, a=4m/s2
distance travelled in time t =1/2*a*(t^2)
distance travelled in 5th half seconds = total distance travelled – distance travelled in 4th half seconds
                                                            = 1/2[a*t*t]-1/2[a*t*t]
                                                             =1/2[4*(5/2)^2]-1/2[4*(4/2)^2]
                                                              =1/2[4*(25-16)/4]
                                                               =1/2*(9)=4.5m
thus, distance travelled in 5th half seconds is 4.5m
 
Harsh Sharma
5 Years ago
the velocity of an object is changing with time and relation is given by the following equation.v=2t+3t^2.calculate the position of the object at t=2s.Assume that the particle to be at origin at t=0.